OpenStudy (anonymous):

solve the equation 1)10t^2=2t^3-72t 2)4(y^2+10)-3y(y+4)=2(y-4)

OpenStudy (solomonzelman):

\(\large\color{black}{ 10t^2=2t^3-72t }\) \(\large\color{black}{ 10t^2 \color{blue}{-10t^2}=2t^3\color{blue}{-10t^2}-72t }\) \(\large\color{black}{ 0=2t^3-10t^2-72t }\)

OpenStudy (solomonzelman):

Factor out of 2t: \(\large\color{black}{ 0=2t(t^2-5t-32) }\)

OpenStudy (solomonzelman):

then you have: \(\large\color{black}{ 0=2t }\) \(\large\color{black}{ 0=t^2-5t-32 }\)